You are the operations manager for an airline and you are considering a higher fare level for passengers in aisle seats. How many randomly selected air passengers must you​ survey? Assume that you want to be 99​% confident that the sample percentage is within 5.5 percentage points of the true population percentage. Complete parts​ (a) and​ (b) below. a. Assume that nothing is known about the percentage of passengers who prefer aisle seats. n= equals 549 ​(Round up to the nearest​ integer.) b. Assume that a prior survey suggests that about 38​% of air passengers prefer an aisle seat. n = equals ?
Answer #1

Solution :

Given that,

(b)

= 0.38

1 - = 0.62

margin of error = E = 0.055

Z/2 = 2.576

sample size = n = (Z / 2 / E)2 * * (1 - )

= (2.576 / 0.055)2 * 0.38 * 0.62

= 517

sample size = n = 517